# A Transfer Attack to Image Watermarks

### v2

@@ Abstract
Watermark has been widely deployed by industry to detect AI-generated images. The robustness of such watermark-based detector against evasion attacks in the white-box and black-box settings is well understood in the literature. However, the robustness in the no-box setting is much less understood. In particular, multiple studies claimed that image watermark is robust in such setting. In this work, we propose a new transfer evasion attack to image watermark in the no-box setting. Our transfer attack adds a perturbation to a watermarked image to evade multiple surrogate watermarking models trained by the attacker itself, and the perturbed watermarked image also evades the target watermarking model. Our major contribution is to show that, both theoretically and empirically, watermark-based AI-generated image detector is not robust to evasion attacks even if the attacker does not have access to the watermarking model nor the detection API.
@@ 1 Introduction
Generative AI (GenAI) can synthesize extremely realistic-looking images, posing growing challenges to information authenticity on the Internet. Watermarking [1, 2, 3, 4, 5, 6, 7] was suggested as a key technology to distinguish AI-generated and non-AI-generated content in the Executive Order on AI security issued by the White House in October 2023. In watermark-based detection, a watermark is embedded into an AI-generated image before releasing it; and an image is detected as AI-generated if the same watermark can be decoded from it. Watermarking AI-generated images has been widely deployed in industry. For instance, Google’s SynthID watermarks images generated by Imagen [8]; OpenAI embeds a watermark into images generated by DALL-E [9]; and Stable Diffusion enables users to embed a watermark into the generated images [10].
@@
An attacker can use evasion attacks [11] to remove the watermark in a watermarked image to evade detection. Specifically, an evasion attack strategically adds a perturbation into a watermarked image such that the target watermark-based detector falsely detects the perturbed image as non-AI-generated. The literature has well understood the robustness of watermark-based detector against evasion attacks in the white-box setting (i.e., the attacker has access to the target watermarking model) and black-box setting (i.e., the attacker has access to the detection API) [11]. Specifically, in the white-box setting, an attacker can find a small perturbation for a given watermarked image such that the perturbed image evades detection while maintaining the image’s visual quality; and in the black-box setting, an attacker can find such a perturbation once the attacker can query the detection API for enough times.
@@
However, robustness of watermark-based detection in the no-box setting (i.e., the attacker does not even have access to the detection API) is much less understood. In such setting, an attacker can use transfer evasion attacks (or transfer attacks for simplicity) [11, 12]. One transfer attack [11] trains one surrogate watermarking model and uses the white-box attack [11] to generate a perturbation for a given watermarked image based on it. An et al. [12], which is concurrent to our work, treat a watermark-based detector as a conventional classifier and propose classifier-based transfer attacks. For instance, given a set of watermarked and non-watermarked images, an attacker can train a surrogate classifier to distinguish them. Given a watermarked image, the attacker uses conventional adversarial example technique [13] to perturb it such that the surrogate classifier predicts the perturbed image as non-watermarked. Prior studies [11, 12] found that existing transfer attacks have limited success at evading advanced watermarking methods [2, 3]. Thus, a common conclusion in prior studies is that watermarking is robust in the no-box setting.
@@
Our work:  In this work, we propose a new transfer attack in the no-box setting to evade watermark-based detection of AI-generated images. Unlike classifier-based transfer attacks [12], our attack directly leverages surrogate watermarking models and thus is better tailored to watermark-based detection. Specifically, an attacker trains multiple surrogate watermarking models using a surrogate dataset. The surrogate watermarking models may have different neural network architectures and watermark lengths, compared to the target watermarking model; and the surrogate dataset may have a different distribution from the one used to train the target watermarking model.
@@
Given multiple surrogate watermarking models, a key challenge is how to find a small perturbation for a given watermarked image such that the perturbed image evades the target watermarking model. We propose a two-step approach to address the challenge. In the first step, the attacker picks a watermark (called target watermark) for each surrogate watermarking model to guide the perturbation search process. Specifically, the perturbation aims to make each surrogate watermarking model decode the target watermark from the perturbed image. For instance, one way is to flip each bit of the watermark decoded by a surrogate watermarking model for the watermarked image and treat the flipped watermark as the corresponding target watermark. Our intuition is that if the perturbation makes multiple surrogate watermarking models decode flipped watermarks from the perturbed image, the target watermarking model is also likely to decode a flipped watermark from the perturbed image, evading detection.
In the second step, given the target watermark for each surrogate watermarking model, we generate a perturbation by aggregating multiple surrogate watermarking models. For instance, one way is to leverage an existing white-box attack [11] to find a perturbation based on each surrogate watermarking model and its target watermark; and then aggregate the multiple perturbations (e.g., take their mean) as the final perturbation. However, such strategy achieves limited success because the aggregation breaks the perturbation patterns, as shown in our experiments. To address the challenge, we find a perturbation by ensembling multiple surrogate watermarking models. Specifically, we formulate an optimization problem, whose objective function is to find a minimum perturbation and constraints are that each surrogate watermarking model decodes its target watermark from the perturbed image. However, the optimization problem is challenging to solve due to the hard constraints. We propose several strategies to reformulate the optimization problem and approximately solve it to find a perturbation.
@@
We theoretically analyze the transferability of our transfer attack. Specifically, we formally quantify the correlation between the target and surrogate watermarking models. Based on such quantified correlation, we derive the probability that the watermark decoded by the target watermarking model from the watermarked image is flipped after adding the perturbation found by our transfer attack to it. Moreover, based on this probability, we further derive both an upper bound and a lower bound for the probability that the watermark decoded by the target watermarking model from the perturbed image matches with the ground-truth watermark in the watermarked image. Such upper and lower bounds quantify the transferability of our attack.
@@
We empirically evaluate our transfer attack on image datasets generated by two GenAI models, i.e., Stable Diffusion and Midjourney. We use HiDDeN [2], which is the state-of-art image watermarking method. Our results show that, with dozens of surrogate watermarking models, our attack successfully evades a watermark-based detector while maintaining image quality, even if the surrogate watermarking models use different neural network architectures and watermark lengths from the target one, and are trained using datasets with different distributions. Moreover, our attack substantially outperforms existing transfer attacks [11, 12]. Our results invalidate the prior beliefs that image watermarks are robust in the no-box setting.
@@
  • We propose a new transfer attack based on multiple surrogate watermarking models to watermark-based AI-generated image detector.
@@
  • We empirically evaluate our transfer attack and compare it with existing ones in different scenarios.
@@ 2.1 Image Watermarks
Three components:  An image watermarking method consists of three components: a watermark that is a bitstring, an encoder that embeds a watermark into an image, and a decoder that decodes a watermark from an image. In non-learning-based methods [5, 1, 14, 15, 16] that have been studied for decades, the encoder and decoder are handcrafted. Unlike these traditional, non-learning-based methods, learning-based methods [2, 4, 3] leverage deep learning. In these methods, both the encoder and decoder are neural networks and trained in an end-to-end manner using an image dataset. The encoder transforms a watermark and an image into feature vectors and combines them to generate a watermarked image; while the decoder outputs a watermark when taking an (unwatermarked or watermarked) image as input. When jointly training the encoder and decoder, the objectives are to minimize the visual difference between an image and its watermarked version while ensuring accurate decoding of the watermark. Compared to non-learning-based ones, learning-based methods are more robust because they can leverage adversarial training [2]. Therefore, we focus on learning-based methods.
Adversarial training:  Adversarial training [13, 17] is a standard method to train robust classifiers and has been extended to train robust watermarking models [2]. The key idea is to add a post-processing layer between the encoder and decoder. The post-processing layer aims to mimic post-processing that a watermarked image may undergo in practice. Specifically, the post-processing layer post-processes a watermarked image before sending it to the decoder. After jointly training the encoder and decoder using adversarial training, the decoder can still decode the watermark in a watermarked image even if it undergoes some post-processing. Thus, we use adversarial training in our experiments to train encoders and decoders.
@@ 2.2 Evasion Attacks
White-box:  Jiang et al. [11] proposed a white-box attack which assumes the attacker has access to the target watermark decoder. Given the target watermark decoder and a watermarked image, an attacker finds a small perturbation such that the watermark decoded from the perturbed image is close to a random watermark. In other words, the perturbation removes the watermark from the watermarked image and thus the perturbed image evades watermark-based detection. Specifically, the attacker finds the perturbation by solving an optimization problem via gradient descent.
@@
Black-box:  In black-box setting, an attacker has access to the API of a target watermark-based detector. Given an image, the detection API returns a binary classification result, i.e., AI-generated (watermarked) or non-AI-generated (non-watermarked). A black-box attack [11] leverages the detection API to perturb a watermarked image to remove its watermark. In particular, given a watermarked image, the attacker starts from an initial perturbed image that is detected as non-AI-generated by the API. The initial perturbed image may have a large perturbation compared to the watermarked image. Then, the attacker repeatedly queries the API and gradually moves the perturbed image towards the watermarked image to reduce the perturbation based on the query results. When an attacker can query the API for enough times, the attacker can find a perturbed image with a small perturbation that evades detection of the API.
No-box:  In this setting, an attacker does not even have access to the detection API. In such setting, common post-processing or transfer attacks can be used to remove watermark from a watermarked image. Specifically, common post-processing refers to common image editing operations such as JPEG compression, Gaussian noise, and Gaussian blur. Watermarks embedded by non-learning-based methods can be removed by common post-processing, but learning-based methods have good robustness against common post-processing due to adversarial training [2]. Existing transfer attacks leverage either only one surrogate watermarking model [11] or classifier-based attacks [12] (these attacks are concurrent to ours). For instance, Jiang et al. [11] proposed to train one surrogate encoder and decoder, and then apply the white-box attack [11] to find a perturbation for a given watermarked image based on the surrogate decoder. An example of classifier-based transfer attack [12] is to train a classifier that distinguishes between watermarked and non-watermarked images; and given a watermarked image, the attacker finds a perturbation such that the perturbed image is classified as non-watermarked by the classifier. As shown in prior works [11, 12] and also confirmed in our experiments, these transfer attacks achieve limited success at evading learning-based watermarking methods.
2.3 Transfer Adversarial Examples
Ensemble-based transfer adversarial examples by leveraging multiple surrogate models have been explored within the adversarial examples community [18, 19]. However, these studies focused on ensemble-based transfer attacks to classifiers, where the surrogate and target models are classifiers. Classifiers are qualitatively different from watermarking decoders. Specifically, a classifier outputs a label for an input, while a watermarking decoder outputs a binary string comprising multiple bits. Such difference results in a fundamentally different optimization objective for transfer attacks. Furthermore, the binary nature of watermark bitstrings enables us to carry out rigorous theoretical analysis on the effectiveness of our transfer attack. However, it is challenging to conduct such theoretical analysis for transfer attacks to classifiers.
@@ 3.1 Watermark-Based Detection
A GenAI service provider trains a watermark encoder (called target encoder) and decoder (called target decoder and denoted as $T$), and uses them for AI-generated image detection. During generation, a watermark $w$ is embedded into each AI-generated image using the target encoder. During detection, a watermark is decoded from a given image $x$ by the target decoder $T$. We denote by $T(x)$ the decoded watermark. We represent the bitwise accuracy between two watermarks $w_{1}$ and $w_{2}$ as $BA(w_{1},w_{2})$, which is the proportion of bits that are identical in $w_{1}$ and $w_{2}$. Watermark-based detection determines whether the image $x$ is AI-generated based on the bitwise accuracy $BA(T(x),w)$ between the decoded watermark $T(x)$ and the ground-truth watermark $w$. Specifically, the image $x$ is detected as AI-generated if the bitwise accuracy $BA(T(x),w)$ either exceeds a certain threshold $\tau$ or falls below $1-\tau$, i.e., $BA(T(x),w)>\tau$ or $BA(T(x),w)<1-\tau$. $\tau$ is commonly set to a value such that the false detection rate, i.e., the probability of falsely detecting a non-AI-generated (i.e., non-watermarked) image as AI-generated (i.e., watermarked), does not exceed a desired small value $\eta$ [11]. For instance, when the watermark has 30 bits and $\eta=10^{-4}$, $\tau\approx 0.83$.
The detector discussed above is known as double-tail detector [11]. In a single-tail detector, the image $x$ is detected as AI-generated if the bitwise accuracy $BA(T(x),w)$ exceeds a threshold $\tau$, i.e., $BA(T(x),w)>\tau$. Double-tail detector is more robust than single-tail detector, e.g., if a perturbed watermarked image evades double-tail detector, then it also evades single-tail detector, but not vice versa. Therefore, we focus on double-tail detector in this work.
@@ 3.2 Threat Model
Attacker’s knowledge:  A GenAI service provider’s watermark-based detector includes a target encoder, a target decoder, a ground-truth watermark $w$, and a detection threshold $\tau$. We assume that the attacker has no-box access to the watermark-based detector. Specifically, the attacker does not have access to the target encoder, target decoder, ground-truth watermark $w$, and detection threshold $\tau$. Moreover, the attacker does not know the neural network architecture of the target encoder and decoder, the length of the ground-truth watermark $w$, nor the image dataset used to train the target encoder/decoder. Such threat model arises when the GenAI service sets its watermark-based detector private and restricts access of the detection API to trusted customers.
@@ 4.1 Overview
We propose a transfer attack to evade watermark-based AI-generated image detection. Our transfer attack trains multiple surrogate watermarking models, and then generates a perturbation for a given watermarked image based on them. We train a set of diverse surrogate watermarking models independently using a dataset that may have a different distribution from the one used by the target watermarking model. For a given watermarked image $x_{w}$, we generate a perturbation $\delta$ based on the surrogate decoders such that the watermark decoded by each surrogate decoder for the perturbed image $x_{w}+\delta$ substantially differs from the watermark decoded for the original watermarked image $x_{w}$. Our intuition is that if multiple surrogate decoders decode a substantially different watermark for the perturbed image $x_{w}+\delta$, it is likely that the target watermarking decoder would decode a substantially different watermark, evading detection. Note that the same surrogate decoders are used to generate perturbations for all watermarked images.
Next, we describe the details of training surrogate models and generating perturbations based on them.

### v4

@@ Abstract
Watermark has been widely deployed by industry to detect AI-generated images. The robustness of such watermark-based detector against evasion attacks in the white-box and black-box settings is well understood in the literature. However, the robustness in the no-box setting is much less understood. In this work, we propose a new transfer evasion attack to image watermark in the no-box setting. Our transfer attack adds a perturbation to a watermarked image to evade multiple surrogate watermarking models trained by the attacker itself, and the perturbed watermarked image also evades the target watermarking model. Our major contribution is to show that, both theoretically and empirically, watermark-based AI-generated image detector based on existing watermarking methods is not robust to evasion attacks even if the attacker does not have access to the watermarking model nor the detection API. Our code is available at: https://github.com/hifi-hyp/Watermark-Transfer-Attack.
@@ 1 Introduction
Generative AI (GenAI) can create highly realistic images, raising concerns about online information authenticity. Watermarking (Bi et al., 2007; Zhu et al., 2018; Tancik et al., 2020; Zhang et al., 2020; Al-Haj, 2007; Abdelnabi & Fritz, 2021; Kirchenbauer et al., 2023) was highlighted as a key technology for distinguishing AI-generated content in the White House’s October 2023 Executive Order on AI security. In this approach, a watermark is embedded in AI-generated images, which can be decoded to verify the image’s origin. Watermarking is widely used, with examples like Google’s SynthID for Imagen (Saharia et al., 2022), OpenAI’s watermark for DALL-E (Ramesh et al., 2021), and Stable Diffusion’s user-enabled watermarking (Rombach, 2022).
@@
An attacker can use evasion attacks (Jiang et al., 2023; Zhao et al., 2023; Saberi et al., 2024; Lukas et al., 2024) to remove a watermark from an image and evade detection. This involves adding a perturbation so that the watermark-based detector falsely identifies the image as non-AI-generated. The robustness of such detectors against evasion attacks has been studied in both white-box (the attacker has access to the watermarking model) and black-box (the attacker can query the detection API) settings (Jiang et al., 2023). In white-box attacks, small perturbations evade detection while preserving image quality, while in black-box attacks, perturbations are found by querying the detection API multiple times.
@@
However, robustness of watermark-based detection in the no-box setting (the attacker even lacks access to the detection API) is less understood. In this setting, attackers may use common post-processing (e.g., JPEG compression) or transfer attacks (Jiang et al., 2023; An et al., 2024). Transfer attacks involve perturbing a watermarked image using surrogate models, which can be classifier-based or watermark-based. Classifier-based attacks treat the detector as a binary classifier and apply adversarial examples (Liu et al., 2017; Chen et al., 2024), while watermark-based attacks train a surrogate watermarking model for white-box attacks (Jiang et al., 2023). Both methods have limited success against advanced watermarking, leading to a misleading conclusion in prior studies that watermarking is robust in the no-box setting.
Figure 1: Watermarked images generated by Stable Diffusion (first row) and their perturbed versions in our transfer attack that successfully evade detection (second row). The target watermarking model uses ResNet architecture. Our transfer attack uses 100 surrogate watermarking models, each of which uses CNN architecture.
@@
Our work:  In this work, we propose a new watermark-based transfer attack in the no-box setting to evade AI-generated image detection. Unlike classifier-based attacks (An et al., 2024), our approach directly uses surrogate watermarking models, making it more suited to watermark-based detection. Unlike previous watermark-based attacks (Jiang et al., 2023), we employ multiple surrogate models instead of one. Specifically, an attacker trains multiple surrogate watermarking models with different architectures and watermark lengths on a surrogate dataset, which may differ in distribution from the one used for the target watermarking model.
@@
A key challenge for our attack is aggregating multiple surrogate watermarking models to find a small perturbation that evades the target watermarking model. We show that simple aggregation methods achieve suboptimal attack effectiveness. To address this, we propose a two-step approach. First, the attacker selects a target watermark for each surrogate model to guide the perturbation search, aiming to make each model decode this target watermark from the perturbed image. For instance, one method flips each bit of the watermark decoded by a surrogate model and uses this as the target. If multiple surrogate models decode flipped watermarks, the target model is likely to do the same, thus evading detection.
In the second step, we generate a perturbation by aggregating multiple surrogate watermarking models. A simple approach is to apply a white-box attack (Jiang et al., 2023) on each model and its target watermark, then average the resulting perturbations. However, this method performs poorly, as aggregation compromises the perturbation patterns. To overcome this, we ensemble the models and formulate an optimization problem that finds a minimum perturbation while ensuring each surrogate model decodes its target watermark. Since this problem is hard to solve, we introduce strategies to reformulate and approximate a solution.
@@
We theoretically analyze the transferability of our attack. First, we quantify the correlation between the target and surrogate watermarking models. Using this correlation, we derive the probability that the target model’s decoded watermark is flipped after adding our perturbation. From this, we further derive upper and lower bounds for the probability that the decoded watermark matches the ground truth. These bounds quantify the transferability of our attack.
@@
We empirically evaluate our transfer attack on image datasets from Stable Diffusion and Midjourney, using multiple watermarking methods (Zhu et al., 2018; Tancik et al., 2020; Fernandez et al., 2023; Jiang et al., 2024). Our attack, using dozens of surrogate models, successfully evades watermark detectors while maintaining image quality (see examples in Figure 1). This holds even when surrogate models differ from the target in algorithms, architectures, watermark lengths, and training datasets. Our attack also outperforms common post-processing, existing transfer attacks (Jiang et al., 2023; An et al., 2024), and the state-of-the-art purification method (Nie et al., 2022), showing that existing image watermarks are broken even in the no-box setting. We note that the effectiveness of our attack to a completely new target watermarking method is unclear, which we discuss in Section 7.
@@
  • We propose a transfer attack based on multiple surrogate watermarking models to watermark-based AI-generated image detector.
@@
  • We empirically evaluate our attack and compare it with existing ones in different scenarios.
@@ 2.1 Image Watermarks
Three components:  An image watermarking method consists of three components: a watermark (a bitstring), an encoder that embeds it into an image, and a decoder that extracts it. In non-learning-based methods (Al-Haj, 2007; Bi et al., 2007; Pereira & Pun, 2000; Kang et al., 2010; Pramila et al., 2018), used for decades, the encoder and decoder are handcrafted but lack robustness to post-processing like JPEG compression or Gaussian noise (Jiang et al., 2023) (as confirmed in Section 7). In contrast, learning-based methods (Zhu et al., 2018; Zhang et al., 2020; Tancik et al., 2020) use deep learning, with both the encoder and decoder trained end-to-end. The encoder combines watermark and image features to generate a watermarked image, while the decoder retrieves the watermark. Joint training minimizes visual differences while ensuring accurate decoding. Learning-based methods, strengthened by adversarial training (Zhu et al., 2018), offer greater robustness, so we focus on them.
Adversarial training:  Adversarial training (Madry et al., 2018; Goodfellow et al., 2014) is a standard method to train robust classifiers and has been extended to train robust watermarking models (Zhu et al., 2018). The key idea is to add a post-processing layer between the watermark encoder and decoder. The post-processing layer aims to mimic post-processing that a watermarked image may undergo. Specifically, the post-processing layer post-processes a watermarked image before sending it to the decoder. After jointly training the encoder and decoder using adversarial training, the decoder can still decode the watermark in a watermarked image even if it undergoes some post-processing. Thus, we use adversarial training in our experiments to train encoders and decoders.
@@ 2.2 Evasion Attacks
White-box:  Jiang et al. (2023) proposed a white-box attack which assumes the attacker has access to the target watermark decoder. Given the target watermark decoder and a watermarked image, an attacker finds a small perturbation such that the watermark decoded from the perturbed image is close to a random watermark. In other words, the perturbation removes the watermark from the watermarked image and thus the perturbed image evades watermark-based detection. Specifically, the attacker finds the perturbation by solving an optimization problem via gradient descent.
@@
Black-box:  In black-box setting, an attacker has access to the API of a watermark-based detector, which detects images as AI-generated (watermarked) or non-AI-generated (non-watermarked). A black-box attack (Jiang et al., 2023; Lukas et al., 2024) uses the API to modify a watermarked image to remove its watermark. Starting with an initially perturbed image detected as non-AI-generated, the attacker gradually reduces the perturbation by repeatedly querying the API. With enough queries, the attacker can find a minimally perturbed image that evades detection.
No-box:  In this setting, an attacker even has no access to the detection API. In such setting, common post-processing or transfer attacks can be used to remove watermarks. Specifically, common post-processing refers to image editing methods such as JPEG compression and Gaussian noise. Watermarks embedded by non-learning-based methods can be removed by common post-processing, but learning-based methods are more robust due to adversarial training (Zhu et al., 2018).
Existing transfer attacks rely on a single surrogate model (Jiang et al., 2023) or conventional classifier-based adversarial examples (An et al., 2024), but achieve limited success against learning-based watermarks (Jiang et al., 2023; An et al., 2024), even if we leverage state-of-the-art multiple surrogate classifiers based transfer adversarial examples (Chen et al., 2024). Note that compared to the existing watermark-based transfer attack (Jiang et al., 2023) that leverages only one surrogate watermarking model, our transfer attack leverages multiple surrogate watermarking models. One technical contribution of our work is to develop methods to aggregate multiple surrogate watermarking models when perturbing a watermarked image. Another technical contribution is that we theoretically analyze the effectiveness of watermark-based transfer attacks.
@@ 3.1 Watermark-Based Detection
A GenAI service provider trains a watermark encoder (target encoder) and decoder (target decoder, denoted $T$) for AI-generated image detection. A watermark $w$ is embedded into each image during generation and decoded from an image $x$ using $T(x)$. The bitwise accuracy $BA(w_{1},w_{2})$ measures the proportion of identical bits between two watermarks $w_{1}$ and $w_{2}$. Detection determines if an image is AI-generated based on the bitwise accuracy $BA(T(x),w)$ between the decoded watermark and the ground-truth watermark. Specifically, $x$ is flagged as AI-generated if $BA(T(x),w)$ exceeds a threshold $\tau$ or falls below $1-\tau$. $\tau$ is set to ensure the false detection rate does not exceed a small value $\eta$ (Jiang et al., 2023) (e.g., for a 30-bit watermark and $\eta=10^{-4}$, $\tau\approx 0.83$).
The above is known as a double-tail detector (Jiang et al., 2023). In a single-tail detector, an image is detected as AI-generated if $BA(T(x),w)>\tau$. The double-tail detector is more robust; if a perturbed watermarked image evades the double-tail detector, it will evade the single-tail detector as well, but not vice versa. Thus, we focus on the double-tail detector in this work.
@@ 3.2 Threat Model
Attacker’s knowledge:  A GenAI service provider’s watermark-based detector includes a target encoder, decoder, ground-truth watermark $w$, and a detection threshold $\tau$. In a no-box setting, the attacker has no access to any of these components, nor the architecture of the encoder/decoder, the length of $w$, or the dataset used for training. This threat model applies when the GenAI service keeps its detector private and limits API access to trusted customers.
@@ 4.1 Overview
We propose a transfer attack to evade watermark-based AI image detection by training multiple surrogate watermarking models. These models, trained independently on a different dataset than the target model, are used to generate perturbations for watermarked images. For a given image $x_{w}$, we generate a perturbation $\delta$ such that the watermark decoded by each surrogate model for $x_{w}+\delta$ differs significantly from the original watermark. The intuition is that if multiple surrogate decoders produce different watermarks, the target decoder will likely do the same, evading detection. The same surrogate decoders are used for all images. Next, we detail the training of surrogate models and perturbation generation.